# Planet-Planet Scattering and ZLK Migration – The Dynamical History of HAT-P-11

### v1

@@ 2.2.2 Four-Planet Simulations
We now consider the effect of planet-planet scattering on the orbit of HAT-P-11 b by running a suite of four-planet scattering simulations. The setup of these simulations is as follows. We perform 10,000 simulations using the new hybrid integrator TRACE (Lu et al., 2024) in REBOUND. We initialize the outer three bodies as in the previous section. However, we now initialize HAT-P-11 b on an $a_{\mathrm{b}}=0.3$ AU orbit (roughly 5×5$\times$5 × its present-day orbit), with all other orbital elements randomized in the same fashion as the three outer planets. The semimajor axis of HAT-P-11 b does not change significantly for most of our simulations. We adopt a timestep of $0.022$ years, which is roughly $1/9$th of HAT-P-11 b’s initialized orbit. The mean fractional energy error over our 10,000 simulations is $10^{-1.35}$.
Figure 4 shows statistics of orbital elements from our simulations. We restrict our discussion to simulations in which HAT-P-11 b survives on an orbit within $a_{\mathrm{b}}<4$ AU; $2449$ simulations satisfy this criterion. Of these simulations, the vast majority ($2111$, or $86\%$) result in three-planet systems comprised of HAT-P-11 b, HAT-P-11 c, and a very distant outer perturber. This is consistent with our results from the previous subsection, which shows that HAT-P-11 b does not play a significant role in the scattering dynamics of the giant planets. $116$ simultions resulted in two-planet systems and $222$ simulations resulted in four-planet systems. We plot the eccentricities and inclinations of all planets corresponding to HAT-P-11 b, HAT-P-11 c and extraneous outer planets in their respective simulations as a function of semimajor axis. The semimajor axis $1\sigma$ distributions are given by $a_{\mathrm{b}}=0.38_{-0.11}^{+0.15}$ AU, $a_{\mathrm{c}}=3.28_{-0.20}^{+0.24}$ AU and $a_{\mathrm{outers}}=44_{-19}^{+110}$ AU. We also see that HAT-P-11 b can be scattered onto a wide range of inclinations and eccentricities. These $1\sigma$ distributions are given by $e_{\mathrm{b}}=0.21^{+0.42}_{-0.17}$ and $\psi_{\mathrm{Ab}}=34_{-21}^{+30}$ degrees.
Figure 4: Results from an ensemble of 10,000 four-planet scattering simulations. We plot the eccentricities and inclinations of the planets representing HAT-P-11 b (green), HAT-P-11 c (maroon) and external planets (blue) as a function of semimajor axis. The distributions of HAT-P-11 c and the outer planets are qualitatively consistent with our results from the three-planet scattering simulations. We see that HAT-P-11 b may be scattered to arbitrarily high eccentricities and very high inclinations.
@@ 3 ZLK Migration
The orbit of HAT-P-11 b is an even bigger puzzle than planet c. Planets as close-in as HAT-P-11 b would naively be expected to have perfectly circular orbits, due to significant tidal effects acting quickly to circularize the orbit. Indeed, the vast majority of hot Jupiters observed do have circular orbits. In addition, its near-polar orbit with $\psi_{Ab}\sim 106^{\circ}$ (Stassun et al., 2017) is also unusual. In this section, we explore ZLK oscillations (driven by HAT-P-11 c) and subsequent tidal circularization and migration as an avenue to explain both of these aspects of HAT-P-11 b’s orbit. We use initial conditions consistent with our four-body scattering simulations.
@@ 3.1 ZLK Oscillations
Lidov (1962) and Kozai (1962) independently discovered a fascinating behavior characteristic of hierarchical three-body systems. Consider such a system, and denote the three bodies as the star, b, and c. If the mutual inclination between the orbits of b and c are significantly misaligned ($39.2^{\circ}<\psi_{\mathrm{bc}}<140.8^{\circ}$), then the orbit of b will experience high-amplitude coupled oscillations in eccentricity and inclination in the absence of short-range forces such as general relativity. Commonly referred to as the Kozai-Lidov effect, the initial discovery of this mechanism by von Zeipel (1910) has been recently brought to light by Ito & Ohtsuka (2019). We hence adopt the name von Zeipel-Lidov-Kozai, or ZLK effect.
For an in-depth review of the ZLK effect as well as its many applications, see Naoz (2016). We provide in this paper a brief overview of the analytic understanding of the ZLK effect necessary to interpret our results. The simplest way to investigate the ZLK effect analytically is in the hierarchical secular approximation, which averages over the mean motion of both b and c (Ford et al., 2000). The validity of this approximation holds as long as $a_{\mathrm{c}}\gg a_{\mathrm{b}}$. The initial parameter space of the HAT-P-11 system in our simulations does not always satisfy this requirement ($a_{\rm b}/a_{\rm c}<0.1$), so we will use direct N-body simulations to explore our system to obtain more accurate results. However, good first-order understanding can still be gained from the secular approximation. In this approximation and assuming b is a test particle, c’s orbit remains unchanged throughout the system’s evolution, and b’s semimajor axis remains fixed. The maximum amplitude of the eccentricity oscillations depends on the initial mutual inclination: assuming the inner begins on a circular orbit, $e_{\mathrm{b,max}}=\sqrt{1-(5/3)\cos^{2}\psi_{\mathrm{bc}}}$ (in the quadrupole limit). The period of the ZLK oscillations is of order (e.g., Fabrycky & Tremaine, 2007):
$\tau_{\rm ZLK}=\frac{2P^{2}_{\rm c}}{3\pi P_{\rm b}}\frac{m_{\mathrm{star}}+m_{\rm b}+m_{\rm c}}{m_{\rm c}}(1-e^{2}_{\rm c})^{3/2}$ (5)
It turns out that the genesis of these effects is the apsidal precession of b’s orbit, $\dot{\omega}_{\rm b}$, generated by the weak perturbations of the outer. In the secular approximation, the apsidal precession rate can be explicitly calculated. The orbit-averaged Hamiltonian of the system is (Ford et al., 2000; Fabrycky & Tremaine, 2007):
@@
This equation can be solved for the apsidal precession rate; see Naoz et al. (2013); Naoz (2016) for an in-depth derivation:
@@
Note that this level of approximation is only strictly valid in the case of a circular outer orbit and a test particle inner, neither of which are good assumptions in our system. The next level of approximation, the octupole-level, would correctly account for both of these complications. However, the dynamics are chaotic in the octupole order (Li et al., 2014). As we aim only to draw first-order analytical understanding from the secular model (to eventually be confirmed with full N-body simulations), we work for now in the quadrupole approximation for greater clarity into the most impactful quantities in this problem.
@@
where $c$ here is the speed of light, not to be confused with a quantity associated with body c. Yee et al. (2018) ruled out the possibility of ZLK oscillations being active in the present-day HAT-P-11 system, on the basis of general relativity suppressing ZLK precession at HAT-P-11 b’s present-day semimajor axis. Plugging in values associated with the present-day system, we see that ZLK oscillations are expected to be quenched by GR precession, with $\dot{\omega}_{\rm b,\mathrm{ZLK}}/\dot{\omega}_{\rm b,\mathrm{GR}}\sim 0.6$. Hence, HAT-P-11 b’s large eccentricity cannot be attributed to it currently being in a high-eccentricity phase of a ZLK oscillation. However, this does not rule out the possibility of ZLK oscillations having occurred in the past. In order for this to have been the case, HAT-P-11 b must have originated at a larger semimajor axis (consistent with our understanding that sub-Neptune-like planets are challenging to form in-situ at the present-day semimajor axis of HAT-P-11 b).
@@ 3.2 Tidal Friction
We turn to the process of ZLK migration, which has been invoked to explain the anomalously high eccentricity and tight orbits of a number of hot Jupiters (see Wu & Murray (2003), Beust et al. (2012) among others). The idea behind ZLK migration is to combine the concepts of ZLK oscillations and tidal evolution. During each high-eccentricity phase of the ZLK oscillation, tidal friction becomes significant due to b’s close approach with the host star.
@@
with $n$ the mean motion of the tidal perturber. We note that more complex and nuanced tidal models are available. In particular, for the high-eccentricity epochs, ZLK oscillations are characterized by dynamic tides may be a more accurate prescription (see Vick et al. (2019) among others). However, no self-consistent N-body package exists with these more complex prescriptions. Hence, we restrict ourselves to the simpler equilibrium tide model in the present work, though we encourage future works to explore the effect of more complex tidal models.
@@
$\begin{split}\dot{\omega}_{\mathrm{Tide}}=&\frac{15\sqrt{G(m_{\mathrm{star}}+m_{\rm b})}}{8a_{\mathrm{b}}^{13/2}}\frac{8+12e_{\mathrm{b}}^{2}+e_{\mathrm{b}}^{4}}{(1-e_{\mathrm{b}}^{2})^{9/2}}\\ &\times\frac{1}{2}\bigg{[}\frac{m_{\rm b}}{m_{\mathrm{star}}}k_{\mathrm{star}}R_{\mathrm{star}}^{5}+\frac{m_{\mathrm{star}}}{m_{\rm b}}k_{\rm b}R_{\rm b}^{5}\bigg{]}.\end{split}$ (11)
Figure 5: Comparison of precessions from the ZLK effect to general relativity and tides in semimajor axis-eccentricity space, using HAT-P-11’s present-day configuration. The maroon line represents where the precessions due to general relativity and ZLK are exactly equal, while the gold lines represent tides with HAT-P-11 b’s present day radius (solid) and tides with twice this radius (dashed). Lines of constant angular momentum are shown in gray. HAT-P-11 b’s present-day position is shown as a black cross. A planet experiencing ZLK-induced high-eccentricity migration will move leftwards in phase space until it encounters one of the colored curves, at which point ZLK oscillations are quenched and it follows one of the contours of constant angular momentum to zero eccentricity.
@@
Figure 5 demonstrates the significance of including the effects of tidal friction. In this figure we consider a system with a central object consistent with HAT-P-11A and an outer perturber consistent with HAT-P-11 c. The semimajor axis and eccentricity of HAT-P-11 b are varied, and we analyze where in semimajor axis-eccentricity space ZLK oscillations are allowed. The present-day orbit of HAT-P-11 b is also marked. The maroon and gold lines mark where $|\dot{\omega}_{\mathrm{ZLK}}/\dot{\omega}_{\mathrm{GR}}|=1$ and $|\dot{\omega}_{\mathrm{ZLK}}/\dot{\omega}_{\mathrm{Tide}}|=1$, respectively. Towards the upper left (lower semimajor axis and higher eccentricity) of each line, ZLK oscillations are quenched by the respective additional effect, while towards the bottom right ZLK oscillations are allowed to be active. From Equation (11), we see that $\dot{\omega}_{\mathrm{Tide}}$ scales strongly with the eccentricity of b. As a result, while tides are less influential than GR at low eccentricities, the opposite is true at higher eccentricities. This is highly impactful for ZLK oscillations: a system where b is at a sufficiently high semimajor axis will be able to sustain ZLK oscillations of extremely high amplitude if only GR is considered, but if tides are accounted for as well during a high-eccentricity epoch the ZLK effect will be quenched.
@@
Consider now the evolution of a body under the influence of tidal effects experiencing ZLK oscillations. If octopule level effects are accounted for (Naoz et al., 2013; Naoz, 2016), the maximum eccentricity of the ZLK oscillations is not constant; they change slowly over the octupole timescale. In addition, during each high-eccentricity epoch tides will shrink the orbit. The result is that over the course of ZLK oscillations, b’s orbit will move towards the upper left corner of Figure 5. Inevitably, at some point there will be a crossing (mostly likely with the gold line), and ZLK oscillations will be quenched. Then b is marooned on a high-eccentricity orbit. This state is only temporary, however, as tides continue to be significant. Now, tidal dissipation acts efficiently to circularize and shrink the planet’s orbit, until b settles into a close-in, circular orbit. Tidal circularization conserves angular momentum $L_{\rm b}$, so b’s evolution through phase space will follow the contour of constant $L_{\rm b}$ (plotted in grey) intersecting the gold line until the planet reaches zero eccentricity.

### v2

@@ 2.2.2 Four-Planet Simulations
We now consider the effect of planet-planet scattering on the orbit of HAT-P-11 b by running a suite of four-planet scattering simulations. The setup of these simulations is as follows. We perform 10,000 simulations using the new hybrid integrator TRACE (Lu et al., 2024) in REBOUND. We initialize the outer three bodies as in the previous section. However, we now initialize HAT-P-11 b on an $a_{\mathrm{b}}=0.3$ AU orbit (roughly 5×5$\times$5 × its present-day semimajor axis), with all other orbital elements randomized in the same fashion as the three outer planets. The semimajor axis of HAT-P-11 b does not change significantly for most of our simulations. We adopt a timestep of $0.022$ years, which is roughly $1/9$th of HAT-P-11 b’s initialized orbit. These simulations are also integrated for $10^{7}$ years. We also extend the maximum simulation distance to $10^{4}$ AU. The mean fractional energy error over our 10,000 simulations is $10^{-1.43}$.
Figure 4 shows statistics of orbital elements from our simulations. We restrict our discussion to simulations in which HAT-P-11 b survives on an orbit within $a_{\mathrm{b}}<4$ AU; $5355$ simulations satisfy this criterion. Of these simulations, the vast majority ($4743$, or $89\%$) result in three-planet systems comprised of HAT-P-11 b, HAT-P-11 c, and a very distant outer perturber. This is consistent with our results from the previous subsection, which shows that HAT-P-11 b does not play a significant role in the scattering dynamics of the giant planets. $611$ simulations resulted in two-planet systems and $1$ simulation resulted in a four-planet system. We plot the eccentricities and inclinations of all planets corresponding to HAT-P-11 b, HAT-P-11 c and extraneous outer planets in their respective simulations as a function of semimajor axis. The semimajor axis $1\sigma$ distributions are given by $a_{\mathrm{b}}=0.29_{-0.06}^{+0.00001}$ AU, $a_{\mathrm{c}}=4.31_{-0.24}^{+0.40}$ AU and $a_{\mathrm{outer}}=59_{-24}^{+78}$ AU. We also see that HAT-P-11 b can be scattered onto a wide range of inclinations and eccentricities. These $1\sigma$ distributions are given by $e_{\mathrm{b}}=0.11^{+0.28}_{-0.07}$ and $\psi_{\mathrm{Ab}}=25_{-17}^{+27}$ degrees. We also ran simulations initializing additional planets between the orbits of HAT-P-11 b and HAT-P-11 c, which are overwhelmingly ejected.
Figure 4: Results from an ensemble of 10,000 four-planet scattering simulations. We plot the eccentricities and inclinations of the planets representing HAT-P-11 b (green), HAT-P-11 c (maroon) and external planets (blue) as a function of semimajor axis. The points referring to HAT-P-11 b and HAT-P-11 c always correspond to the innermost planet and the innermost giant planet in our simulations, respectively. In every simulation these are distinct. The distributions of HAT-P-11 c and the outer planets are qualitatively consistent with our results from the three-planet scattering simulations. We see that HAT-P-11 b may be scattered to arbitrarily high eccentricities and very high inclinations.
@@ 3 ZLK Migration
The orbit of HAT-P-11 b is an even bigger puzzle than planet c. Its near-polar orbit with $\psi_{Ab}\sim 106^{\circ}$ (Stassun et al., 2017) is highly unusual. Most intriguingly, planets as close-in as HAT-P-11 b would naively be expected to have perfectly circular orbits, due to significant tidal effects acting quickly to circularize the orbit. Indeed, the vast majority of hot Jupiters observed do have circular orbits (Dawson & Johnson, 2018). To first order, the tidal circularization timescale is given by (Goldreich & Soter, 1966):
$t_{\mathrm{circ}}=\frac{4}{63}\frac{a^{13/2}}{\sqrt{GM_{*}^{3}}}Q_{p}m_{p}R_{p}^{-5}.$ (5)
Plugging in the present-day values of the system and $Q_{p}=10^{5}$, a fiducial estimate for the tidal quality factor of a Neptune-like planet (e.g Millholland, 2019), we obtain $t_{\mathrm{circ}}=2\times 10^{9}$ years, which is compatible with the system’s age of $6.5^{+5.9}_{-4.1}$ Gyr (Yee et al., 2018). However, these expressions must be used with caution. Wisdom (2008) found that the standard analytic expressions can underestimate tidal dissipation by several orders of magnitude for high eccentricities, and expressions such as Equation (5) significantly overpredict the circularization timescale. Given this, and the fact that tides act more efficiently at the highest eccentricities, we naively expect that HAT-P-11 b should be circularized.
However, we cannot state with certainty that the orbit should be circularized. First, we note that the default calculation of $t_{\mathrm{circ}}$ is not that far off from the system’s age. In addition, the unconstrained tidal quality factor $Q_{p}$ may well be higher than the nominal value of $10^{5}$ as argued by Mardling (2008). In this case, HAT-P-11 b could have formed at its present-day semimajor axis and gained significant eccentricity/spin-orbit misalignment through scattering alone, without having sufficient time to fully circularize. Given the timescales and assumptions enumerated, we consider this scenario a somewhat unlikely but certainly possible scenario.
In this section, we explore a more likely alternative, where ZLK oscillations (driven by HAT-P-11 c) and subsequent tidal circularization and migration act as an avenue to explain both the eccentricity and spin-orbit misalignment of HAT-P-11 b. We use initial conditions consistent with our four-body scattering simulations.
@@ 3.1 ZLK Oscillations
Lidov (1962) and Kozai (1962) independently identified a fascinating behavior characteristic of hierarchical three-body systems. Consider such a system, and denote the three bodies as the star, b, and c. If the mutual inclination between the orbits of b and c are significantly misaligned ($39.2^{\circ}<\psi_{\mathrm{bc}}<140.8^{\circ}$), then the orbit of b will experience high-amplitude coupled oscillations in eccentricity and inclination in the absence of short-range forces such as general relativity. Commonly referred to as the Kozai-Lidov effect, the initial discovery of this mechanism by von Zeipel (1910) has been recently brought to light by Ito & Ohtsuka (2019). We hence adopt the name von Zeipel-Lidov-Kozai, or ZLK effect.
For an in-depth review of the ZLK effect as well as its many applications, see Naoz (2016). We provide in this paper a brief overview of the analytic understanding of the ZLK effect necessary to interpret our results. The simplest way to investigate the ZLK effect analytically is in the hierarchical secular approximation, which averages over the mean motion of both b and c (Ford et al., 2000). The validity of this approximation holds as long as $a_{\mathrm{c}}\gg a_{\mathrm{b}}$. The initial parameter space of the HAT-P-11 system in our simulations does not always satisfy this requirement ($a_{\rm b}/a_{\rm c}<0.1$), so we will use direct N-body simulations to explore our system to obtain more accurate results. However, good first-order understanding can still be gained from the secular approximation [^1] The extent to which the secular approximation is valid for mildly hierarchical systems is discussed in Grishin et al. (2018).. In this approximation and assuming b is a test particle, c’s orbit remains unchanged throughout the system’s evolution, and b’s semimajor axis remains fixed. The maximum amplitude of the eccentricity oscillations depends on the initial mutual inclination: assuming the inner begins on a circular orbit, $e_{\mathrm{b,max}}=\sqrt{1-(5/3)\cos^{2}\psi_{\mathrm{bc}}}$ (in the quadrupole limit). The period of the ZLK oscillations is of order (e.g., Fabrycky & Tremaine, 2007):
$\tau_{\rm ZLK}=\frac{2P^{2}_{\rm c}}{3\pi P_{\rm b}}\frac{m_{\mathrm{star}}+m_{\rm b}+m_{\rm c}}{m_{\rm c}}(1-e^{2}_{\rm c})^{3/2},$ (6)
where $P_{\mathrm{b}},P_{\mathrm{c}}$ are the periods of the inner and outer bodies, respectively.
The genesis of these effects is the apsidal precession of b’s orbit, $\dot{\omega}_{\rm b}$, generated by the weak perturbations of the outer. In the secular approximation, the apsidal precession rate can be explicitly calculated. The orbit-averaged Hamiltonian of the system is (Ford et al., 2000; Fabrycky & Tremaine, 2007):
@@
where $\psi_{\mathrm{bc}}$ is the mutual inclination between bodies $b$ and $c$. This equation can be solved for the apsidal precession rate; see Naoz et al. (2013); Naoz (2016) for an in-depth derivation:
@@
where $G$ is the gravitational constant, $L_{\mathrm{b/c}}$ are the Delaunay action variables, and $G_{\mathrm{c}}\equiv L_{\mathrm{c}}\sqrt{1-e_{\mathrm{c}}^{2}}$. Note that this level of approximation is only strictly valid in the case of a circular outer orbit and an inner test particle, neither of which are good assumptions in our system. The next level of approximation, the octupole-level, would correctly account for both of these complications. However, the dynamics are chaotic in the octupole order (Li et al., 2014a). As we aim only to draw first-order analytical understanding from the secular model (to eventually be confirmed with full N-body simulations), we work for now in the quadrupole approximation for greater clarity into the most impactful quantities in this problem.
@@
where $c$ here is the speed of light, not to be confused with a quantity associated with body c. Yee et al. (2018) ruled out the possibility of ZLK oscillations being active in the present-day HAT-P-11 system, on the basis of general relativity suppressing ZLK precession at HAT-P-11 b’s present-day semimajor axis. Plugging in values associated with the present-day system, we confirm that ZLK oscillations are expected to be quenched by GR precession, with $\dot{\omega}_{\rm b,\mathrm{ZLK}}/\dot{\omega}_{\rm b,\mathrm{GR}}\sim 0.8$. Hence, HAT-P-11 b’s large eccentricity cannot be attributed to it currently being in a high-eccentricity phase of a ZLK oscillation. However, this does not rule out the possibility of ZLK oscillations having occurred in the past. In order for this to have been the case, HAT-P-11 b must have originated at a larger semimajor axis.
@@ 3.2 Tidal Friction
We turn to the process of ZLK migration, which has been invoked to explain the anomalously high eccentricity and tight orbits of a number of hot Jupiters (see Wu & Murray (2003), Mardling (2010), Beust et al. (2012) among others). The idea behind ZLK migration is to combine the concepts of ZLK oscillations and tidal evolution. During each high-eccentricity phase of the ZLK oscillation, tidal friction becomes significant due to b’s close approach with the host star.
@@
with $n$ the mean motion of the tidal perturber, in this case the host star. We note that more complex and nuanced tidal models are available. In particular, for the high-eccentricity ZLK epochs dynamical tides may be a more accurate prescription (e.g Mardling, 1995; Lai, 2012; Fuller et al., 2014; Vick et al., 2019, 2023). However, no self-consistent N-body package exists with these more complex prescriptions. Hence, we restrict ourselves to the simpler equilibrium tide model in the present work, though we encourage future works to explore the effect of more complex tidal models.
@@
$\begin{split}\dot{\omega}_{\mathrm{Tide}}=&\frac{15\sqrt{G(m_{\mathrm{star}}+m_{\rm b})}}{8a_{\mathrm{b}}^{13/2}}\frac{8+12e_{\mathrm{b}}^{2}+e_{\mathrm{b}}^{4}}{(1-e_{\mathrm{b}}^{2})^{9/2}}\\ &\times\frac{1}{2}\bigg{[}\frac{m_{\rm b}}{m_{\mathrm{star}}}k_{\mathrm{star}}R_{\mathrm{star}}^{5}+\frac{m_{\mathrm{star}}}{m_{\rm b}}k_{\rm b}R_{\rm b}^{5}\bigg{]},\end{split}$ (12)
@@
where $k_{\mathrm{star}},k_{\mathrm{b}}$ are the tidal Love numbers of the star and planet, respectively. Figure 5 demonstrates the significance of including the effects of tidal friction. In this figure we consider a system with a central object consistent with HAT-P-11A and an outer perturber consistent with HAT-P-11 c. The semimajor axis and eccentricity of HAT-P-11 b are varied, and we analyze where in semimajor axis-eccentricity space ZLK oscillations are allowed. The present-day orbit of HAT-P-11 b is also marked. The maroon and gold lines mark where $|\dot{\omega}_{\mathrm{ZLK}}/\dot{\omega}_{\mathrm{GR}}|=1$ and $|\dot{\omega}_{\mathrm{ZLK}}/\dot{\omega}_{\mathrm{Tide}}|=1$, respectively. Towards the upper left (lower semimajor axis and higher eccentricity) of each line, ZLK oscillations are quenched by the respective additional effect, while towards the bottom right ZLK oscillations are allowed to be active. From Equation (12), we see that $\dot{\omega}_{\mathrm{Tide}}$ scales strongly with the eccentricity of b. As a result, while tides are less influential than GR at low eccentricities, the opposite is true at higher eccentricities. This is highly impactful for ZLK oscillations: a system where b is at a sufficiently high semimajor axis will be able to sustain ZLK oscillations of extremely high amplitude if only GR is considered, but if tides are accounted for as well during a high-eccentricity epoch the ZLK effect will be quenched.
Figure 5: Comparison of precessions from the ZLK effect to general relativity and tides in semimajor axis-eccentricity space, using HAT-P-11’s present-day configuration. The maroon line represents where the precessions due to general relativity and ZLK are exactly equal, while the gold lines represent tides with HAT-P-11 b’s present day radius (solid) and tides with twice this radius (dashed). Lines of constant angular momentum are shown in gray. HAT-P-11 b’s present-day position is shown as a black cross. A planet experiencing ZLK-induced high-eccentricity migration will move towards the upper left hand corner in eccentricity-semimajor axis space until it encounters one of the colored curves, at which point ZLK oscillations are quenched and it follows one of the contours of constant angular momentum to zero eccentricity.
@@
Consider now the evolution of a body under the influence of tidal effects experiencing ZLK oscillations. If octupole level effects are accounted for (Naoz et al., 2013; Naoz, 2016), the maximum eccentricity of the ZLK oscillations is not constant; rather, the maxima increase slowly over the octupole timescale (Naoz et al., 2013; Li et al., 2014b). In addition, during each high-eccentricity epoch tides will shrink the orbit. The result is that over the course of ZLK oscillations, b’s orbit will move towards the upper left corner of Figure 5. Inevitably, at some point there will be a crossing (mostly likely with the gold line), and ZLK oscillations will be quenched. Then b is marooned on a high-eccentricity orbit. This state is only temporary, however, as tides continue to be significant. Now, tidal dissipation acts efficiently to circularize and shrink the planet’s orbit, until b settles into a close-in, circular orbit. Tidal circularization conserves angular momentum $L_{\rm b}$, so b’s evolution through phase space will follow the contour of constant $L_{\rm b}$ (plotted in grey) intersecting the gold line until the planet reaches zero eccentricity.